The length of a rectangle is 5 yards more than the width! If the perimeter is 42 yards, what are the length and width?
Answer by macston(5194)

W=width; L=length-W+5yds; P=Perimeter=2(L+W)=42 yds
P=2(L+W) Substitute for L.
42yds=2((W+5yds)+W) Divide each side by 2.
21yds=2W+5yds Subtract 5yds from each side.
16yds=2W Divide each side by 2.
8yds=W ANSWER 1: The width is 8 yds.
L=W5 yds=8 yds+5 yds=13 yds ANSWER 2: The length is 13 yds.
CHECK:
P=2(L+W)
42 yds=2(13 yds + 8 yds)
42 yds=2(21 yds)
42 yds=42 yds